(a) Transform the given equation xy=a in Polar coordinates
1
Expert's answer
2021-08-16T11:20:53-0400

Let's move on to polar coordinates

"x = r\cos \varphi ,\,y = r\sin \varphi"

We get the equation

"r\cos \varphi \cdot r\sin \varphi = a \Rightarrow {r^2} \cdot \frac{1}{2}\sin 2\varphi = a \Rightarrow {r^2} = \frac{{2a}}{{\sin 2\varphi }}"

Answer: "{r^2} = \frac{{2a}}{{\sin 2\varphi }}"